[figure: 9.7 cm]

THE MAGNETIC PROPERTIES AND THE BARKHAUSEN NOISE OF THE HYDROGENED Fe-V-B AMORPHOUS ALLOY

L. Ceniga^{a}, L. Novák^{b}

aInstitute of Experimental Physics SAS, Watsonova 47, 043 53 Košice, Slovak Republic

bDepartment of Physics, Technical University, Komenského park 2, 041 20 Košice, Slovak Republic

In order to get information about the influence of the hydrogenation-dehydrogenation process, the Fe-V-B amorphous alloy was hydrogenated by an electrolytical method. Barkhausen noise parameters and magnetic properties as the total anisotropy, the stress induced anisotropy, the coercive force, the remanent magnetization, the saturation magnetization, the demagnetizing factor were measured before hydrogenation and after it as a function of time at room temperature. Inner stress centres around Bernal’s holes ocuppied by hydrogen atoms were responsible for rapid changes of the magnetic properties and the Barkhausen noise parameters. The hydrogenation-dehydrogenation process caused structural relaxation and destroyed as-cast clusters.

[figure: 13.4 cm]

Subject category :

3. Transition Metals, Alloys and Compounds

Presentation mode : poster

Corresponding author : L. Ceniga

Address for correspondence : Institute of Experimental Physics SAS Watsonova 47 043 53 Košice Slovak Republic

Email address : ceniga@saske.sk